Intervento di GIANCARLO FRANZESE

Titolo:

The fully frustrated XY model with next nearest neighbor interaction

Autori: Giancarlo Franzese(1,4), Vittorio Cataudella(2,4), Rosario Fazio(3,4)

(1) Dip. di Fisica “E.Amaldi”, Università Roma Tre, via della Vasca Navale 84, 00146 Roma - Italy (Present address: Center for Polymer Studies, Boston University)

(2) Dip. di Scienze Fisiche, Universitá “Federico II”, Mostra d’Oltremare, Pad. 19, 80125 Napoli - Italy

(3) Dip. di Metodologie Chimiche e Fisiche per l’Ingegneria (DMCFI), Uni. di Catania, viale A. Doria 6, 95125 Catania - Italy

(4) Istituto Nazionale per la Fisica della Materia (INFM), Unitá di Napoli e Catania

We introduce [1] a fully frustrated XY model with nearest neighbor (nn) and next nearest neighbor (nnn) couplings which could play a relevant role in low temperature regime of usual Josephson junction arrays and can be realized in specially designed setups [2]. We study the phase diagram for 0 ≤ x ≤ 1 (x is the ratio between nn and nnn couplings). When x ≤ 0.7 an Ising-like and a Kosterlitz-Thouless-Berezinskii-like transitions are present (both decrease with increasing x). For x ≥ 0.8 only the latter transition occurs. Above x = 0.8 the transition temperature rapidly increases toward the x → ∞ limit and the reentrant phase could be experimentally detected.

Referenze:

[1] V. Cataudella, G. Franzese, R. Fazio, in preparation.

[2] H. Courtois and P. Martinoli (private communication).